# MuLBA: Multidimensional Adaptive Language-based Assessments for Overlapping Mental Health Conditions

### v2

@@ Abstract
Recent advances in large language models (LLMs) offer new opportunities for scalable, interactive mental health assessment, but excessive querying by LLMs burdens users and is inefficient for real-world screening across transdiagnostic symptom profiles. We introduce MAQuA, an adaptive question-asking framework for simultaneous, multidimensional mental health screening. Combining multi-outcome modeling on language responses with item response theory (IRT) and factor analysis, MAQuA selects the questions with most informative responses across multiple dimensions at each turn to optimize diagnostic information, improving accuracy and potentially reducing response burden. Empirical results on a novel dataset reveal that MAQuA reduces the number of assessment questions required for score stabilization by 50–87% compared to random ordering (e.g., achieving stable depression scores with 71% fewer questions and eating disorder scores with 85% fewer questions). MAQuA demonstrates robust performance across both internalizing (depression, anxiety) and externalizing (substance use, eating disorder) domains, with early stopping strategies further reducing patient time and burden. These findings position MAQuA as a powerful and efficient tool for scalable, nuanced, and interactive mental health screening, advancing the integration of LLM-based agents into real-world clinical workflows.
Introduction
@@
Recent progress in large language models (LLMs) has enabled the automated inference of mental health scores from patient-generated natural language. However, comprehensive evaluations indicate that such LLM-based assessments are inconsistent (Ji et al. 2022) and, in many cases, less accurate than dedicated, condition-specific models with established psychometric validity (Harrigian, Aguirre, and Dredze 2020). These limitations present critical barriers to the clinical adoption and trustworthiness of LLMs in the mental health domain.
@@
Traditional NLP approaches in mental health have often relied on annotations of specific conditions derived from social media, primarily focusing on single-task models, such as encoder-based classifiers for depression (Coppersmith et al. 2015; Eichstaedt et al. 2018), suicidal ideation (Shen et al. 2017) and anxiety (Owen, Camacho-Collados, and Espinosa Anke 2020; Gkotsis et al. 2017), detection. The narrow scope of modeling a single dimensional mental health score typically fails to capture comorbidities or the complex, multidimensional nature of symptoms observed by real-world clinicians (Shani and Stade 2025). More importantly, they do not address the interactive paradigm where LLM agents engage with users in prompted settings: language generated in response to structured diagnostic interviews, as would be typical in real-world clinical settings.
@@
Further, in actual clinical practice, clinicians dynamically adapt their lines of questioning based on prior information received, avoiding redundancy, clarifying ambiguous responses, and addressing emergent concerns (James, Morse, and Howarth 2010). While LLMs excel at modeling linguistic patterns, they often struggle to dynamically ground inferred states in underlying mental health constructs (Singh et al. 2025), especially given the multi-objective challenge of simultaneously selecting diagnostic questions and evaluating mental health status (Li et al. 2025; Sener and Koltun 2018). Furthermore, maximizing the diagnostic yield within the constraints of limited clinician-patient interaction time remains a key priority, as excessive probing, especially via LLM-based dialogue agents can be mentally taxing and lead to decision fatigue or disengagement (Jin, Kim, and Han 2025), highlighting the need for adaptive systems that select only the most informative and relevant follow-up questions.
@@
To address these challenges, we propose MAQuA: an adaptive, language-based assessment framework that supports multidimensional mental health modeling. This framework infers multiple underlying condition scores while adaptively selecting the most informative follow-up questions, thereby guiding interactions efficiently toward richer diagnostic insight, making it suitable to operate alongside LLM agents. Building on item response theory (IRT)-based adaptive assessments introduced by  (Varadarajan et al. 2024), our results demonstrate that optimizing information gain across multiple conditions simultaneously can be even more effective than single-score models.
@@
To explore this, we empirically benchmark single-task and multitask models on a multidimensional mental health dataset and assess the effectiveness of adaptive question selection. Our research investigates (1) whether cross-condition information sharing improves per-condition predictive performance; (2) construct validity of automatically inferred mental health dimensions; and (3) the efficacy of multidimensional IRT in sustaining validity of measures in subsequent adaptive assessment turns. Our main contributions include: (a) a systematic comparison of multi-condition modeling techniques, (b) an adaptive assessment framework for optimizing information gain across outcomes, (c) empirical results demonstrating robust gains in multidimensional assessment, and (d) the release of a novel, questionnaire-driven dataset to support future research.
Related Work
@@
While large language models (LLMs) have shown promising results in zero-shot and few-shot prediction tasks (Ganesan et al. 2024; Hur et al. 2024), finetuned or instruction-tuned models remain generally more reliable and better validated across a range of mental health outcomes (Xu et al. 2024). Recent advances demonstrate that open-ended language responses to standardized questions can predict mental health scores with high accuracies: sometimes with correlations exceeding 0.8 with established clinical rating scales (Kjell et al. 2022; Varadarajan et al. 2024; Sikström et al. 2023). These include pre-trained language models tailored for mental healthcare applications, such as ClinicalBERT and MentalBERT (Alsentzer et al. 2019; Ji et al. 2022).
@@
Large language models (LLMs) commonly address referential and vague queries by employing targeted, selective prompts, which has been shown to improve answer accuracy and reduce errors (Zhang, Knox, and Choi 2025; Kuhn, Gal, and Farquhar 2023). However, despite these improvements, LLMs still fall short of human conversational subtlety and adaptiveness when it comes to clarification and follow-up questions. In terms of LLMs for mental health support, Rosenman, Hendler, and Wolf (2024) demonstrate that LLMs can effectively transform unstructured psychological interviews into structured questionnaires, enabling automated, multidimensional psychiatric evaluation, though reliability and consistency still require further improvement for clinical deployment. Complementing this, Nguyen et al. (2025) explore LLMs’ ability to engage in mental health counseling, showing that, though LLMs can generate contextually relevant follow-up questions, they often lag behind human clinicians in empathy, specificity, and diagnostic nuance, and in crafting clarifying or probing questions that are crucial for effective counseling. Similarly, Yang et al. (2023) assess LLM performance across a spectrum of mental health tasks, finding that LLMs frequently overlook emotional cues or oversimplify questions, limiting their utility for nuanced clinical interpretation. This underscores the need for methods like ours that explicitly model multiple mental health factors to guide strategic question selection. Our framework enhances both the efficiency and accuracy of mental health assessments by optimizing inquiry and serving as a comprehensive diagnostic tool.
@@
While adaptive testing has gained significant traction in educational settings, its multidimensional applications remain relatively underexplored, particularly when leveraging open-ended language responses. Most existing approaches in NLP rely on unidimensional item response theory (IRT) focusing on single outcomes (Lalor, Wu, and Yu 2016; Varadarajan et al. 2024). To date, no prior work has effectively bridged this gap by integrating adaptive item selection with multitask learning for language-based, multidimensional mental health evaluation.
@@
Item Response Theory (IRT) is a probabilistic, data-driven measurement framework that models the relationship between an individual’s latent trait score (such as depression severity) and their probability of specific item responses on a questionnaire (Embretson and Reise 2000; Hambleton, Swaminathan, and Rogers 1991). In single-dimensional IRT, this relationship is modeled with respect to one latent trait (usually denoted $\theta$), accounting for item-specific parameters such as difficulty and discrimination. IRT enables precise ordering and calibration of items based on their informativeness in measuring the latent trait, supporting adaptive and individualized assessment. Adaptive language-based assessments (Varadarajan et al. 2024, ALBA) were first introduced using single-dimensional IRT, summarized in Algorithm 1.
Exploratory Factor analysis (EFA) is a statistical technique used to uncover the underlying structure of a set of observed variables by identifying clusters of variables that co-vary together, known as factors or latent constructs (Cudeck and MacCallum 2000). Although it is very similar to Principal Component Analysis (PCA), it captures the multidimensional, often overlapping nature of latent variables by allowing factors to be correlated, unlike PCA which assumes uncorrelated components. This approach captures meaningful psychological constructs like depression and anxiety, supporting the development of sensitive, multidimensional assessments for accurate mental health screening. EFA allows us to distill large and complex data such as responses to multiple questionnaires assessing various psychological conditions. Each factor represents a distinct underlying construct that accounts for shared variance among the observed variables, providing insight into how different mental health symptoms or traits may be related at a deeper level.
@@
Dataset
To explore these questions, we collected a unique dataset of language questions, diagnoses, validated clinical scales to administer. We collected the dataset in two phases: first, the participants were pre-screened to establish a diverse sampling pool, after which mental health data were collected from these screened individuals using standardized rating scales, free-text narratives, and open-response questions targeting nine common mental disorders, including mood disorders (major depression, generalized anxiety disorder, bipolar disorder), autism spectrum disorder, attention-deficit/hyperactivity disorder, eating disorders, obsessive-compulsive disorder, post-traumatic stress disorder, and substance use disorders (alcohol and/or drug abuse).
We recruited a small set of 523 participants (1041 recruited, 523 completed) who were diagnosed with any of nine mental health conditions (Anxiety Disorders (AD), Bipolar Disorder (BD), Depression (D), Attention Deficit Hyperactivity Disorder (ADHD), Post-Traumatic Stress Disorder (PTSD), Obsessive-Compulsive Disorder (OCD), Eating Disorders (ED), Addiction and Substance Abuse (A), Autism (AU)). We screened participants for having one (or more) ongoing mental disorders (e.g., Np = 523, $\sim$50 of each disorder). The distribution of all the diagnoses across the participants is shown in Figure A.1. The participants first took a screening questionnaire for diagnoses and treatment for the ten mental health conditions to qualify for eligibility to participate in the survey.
@@
Ten validated clinical scales were also administered: the PHQ-9 ( Kroenke, Spitzer, and Williams (2001); 9 items, 4-point Likert) for depression, GAD-7 (Spitzer et al. (2006); 7 items, 4-point Likert) for anxiety, MDQ (Miller et al. (2004); 14 binary items plus one 4-point Likert item) for bipolar disorder, RAADS-14 (Eriksson, Andersen, and Bejerot (2013); 14 items, 4-point Likert) for autism, ASRS Part A (Adler et al. (2006); 6 items, 5-point Likert) for ADHD, NSESSS-PTSD (LeBeau et al. (2014); 9 items, 5-point Likert plus an open-text trauma description) for PTSD, BOCS (Goodman et al. (1989); 15 items, 3-point Likert plus an open-response categorization) for obsessive-compulsive symptoms, EDE-QS (Fairburn and Beglin (2008); 12 items, 4-point Likert) for eating disorders, and two substance use instruments: the AUDIT (Allen et al. (1997); 8 items, 5-point Likert plus 2 items, 3-point Likert) for alcohol misuse and the DUDIT (Berman et al. (2007); 9 items, 5-point Likert plus 2 items, 3-point Likert) for drug abuse. Post-screening, we invited the selected participants to participate in the main study, with recruitment conducted via Prolific. The language questions, some eliciting descriptive words and some open-ended essay-like responses, are listed in Table A.2. They were asked in random order to eliminate any systematic priming effects.
Methods
@@
7:  Set initial latent trait estimates $\boldsymbol{\theta}_{i}^{(0)}$ for each individual $i$
@@
9:while stopping criteria not met on $\boldsymbol{\theta}_{i}^{(t)}$ and pool not empty do
10:  For each candidate item $p$, compute Fisher information matrix $\mathcal{I}_{p}(\boldsymbol{\theta}_{i}^{(t)})$
11:  Select next item $p^{*}$ maximizing $\det(\mathcal{I}_{p^{*}}(\boldsymbol{\theta}_{i}^{(t)}))$ over all candidates (D-optimality)
@@
15:  Update $\boldsymbol{\theta}_{i}^{(t+1)}$ using maximum likelihood on responses
@@
18:Output: Final multidimensional trait estimates $\boldsymbol{\theta}_{i}^{(t)}$ and diagnostic profile for individual $i$
Multi-outcome Modeling
Given the multitude of psychological dimensions, associated language data, and the occurrence of comorbid diagnoses, linguistic expressions intended to capture one dimension may also provide valuable information about others. To investigate this, we frame the language-based modeling of psychological scores in two configurations: single-task and multi-task. Using a stratified sampling approach based on depression outcomes (PHQ scores) over the users, we generate 9 folds for evaluation, 5 as train set, 3 as development set and 1 test set. Each fold contains the responses of around 58 users each. Language representations are derived from all text responses utilizing the nomic-embed-text-v1.5 model (Nussbaum et al. 2024), which are subsequently reduced to 16 dimensions through Matryoshka embeddings (Kusupati et al. 2022). We then train linear regression models to predict standard mental health questionnaire scores across ten dimensions (including PHQ for depression, GAD for anxiety etc. as explained in sec2). The single-task models predict each mental health dimension independently, whereas the multi-task models simultaneously predict all ten scores (see Figure 1). For model optimization, we performed hyperparameter tuning over ranges of learning rates, weight decay values, and used the Adam optimizer. We also compared output scaling methods for the regression, finding that min-max scaling consistently outperformed z-score normalization across all settings. Experiments were conducted using a single NVIDIA A6000 GPU. Further, we explore several model variants to analyze different aspects of the language data.
@@
Valid. Scales
- .404 .423 .440 .454 .097 .133 .073 .172 .080 - .253
@@
1. Aggregation type
2. Question Information
Factor Analysis and MIRT
Multidimensional adaptive question-asking algorithm requires setting up a factor model, which is defined through factor analysis to determine the factor structure, that takes in all the questions.
Figure 2: Question texts loading on to the two factors.
We run exploratory factor analysis to determine the optimal factor structure for this dataset for the predicted user-level scores. The factor loadings are reported in Table 3 and Figure 2. Then we run aggregate the multi-outcome model predictions at a question level, across all users, to analyze how each question loads on to the factors by by applying the factor model at the question-level. This yields a question-level loading which is used to define the MIRT model (See Table A.3 for details.)

### v3

# MAQuA: Multi-outcome Adaptive Question-Asking for Mental Health using Item Response Theory

@@ Abstract
Recent advances in LLMs offer new opportunities for scalable, interactive mental health assessment, but excessive querying burdens users and is inefficient for real-world assessments across transdiagnostic symptom profiles. We introduce MAQuA, a multi-outcome modeling and adaptive question-asking framework for simultaneous, multidimensional mental health assessment. Combining multi-outcome modeling on language responses with item response theory (IRT) and factor analysis, MAQuA selects the questions with most informative responses across multiple dimensions at each turn to optimize diagnostic information, improving accuracy and potentially reducing response burden. Empirical results on a novel dataset reveal that MAQuA reduces the number of assessment questions required for score stabilization by 50–87% compared to random ordering (e.g., achieving stable depression scores with 71% fewer questions and eating disorder scores with 85% fewer questions). MAQuA demonstrates robust performance across both internalizing (depression, anxiety) and externalizing (substance use, eating disorder) domains, with early stopping strategies further reducing patient time and burden. These findings position MAQuA as a powerful and efficient tool for scalable, nuanced, and interactive mental health assessment, advancing the integration of LLM-based agents into real-world clinical workflows.
MAQuA: Multi-outcome Adaptive Question-Asking for Mental Health using Item Response Theory
[…]
1 Introduction
@@
Recent progress in large language models (LLMs) has enabled the automated inference of mental health scores from patient-generated natural language. However, comprehensive evaluations indicate that such LLM-based assessments are inconsistent ji2022mentalbert and, in many cases, less accurate than dedicated, condition-specific models with established psychometric validity harrigian2020models. These limitations present critical barriers to the clinical adoption and trustworthiness of LLMs in the mental health domain.
@@
Traditional NLP approaches in mental health have often relied on annotations of specific conditions derived from social media, primarily focusing on single-task models, such as encoder-based classifiers for depression coppersmith2015clpsych; eichstaedt2018facebook, suicidal ideation or risk shen2017depression; varadarajan-etal-2024-archetypes and anxiety owen2020towards; gkotsis2017characterisation; juhng2023discourse, detection. The narrow scope of modeling a single dimensional mental health score typically fails to capture comorbidities or the complex, multidimensional nature of symptoms observed by real-world clinicians shani-stade-2025-measuring; soni-etal-2025-mental; varadarajan-etal-2025-linking. More importantly, they do not address the interactive paradigm where LLM agents engage with users in prompted settings: language generated in response to structured diagnostic interviews, as would be typical in real-world clinical settings.
@@
Further, in actual clinical practice, clinicians dynamically adapt their lines of questioning based on prior information received, avoiding redundancy, clarifying ambiguous responses, and addressing emergent concerns james2010science; welch2025isca. While LLMs excel at modeling linguistic patterns, there are mixed signals on their ability to dynamically ground inferred states in underlying mental health constructs singh-etal-2025-systematic; ganesan2024explaining, especially given the multi-objective challenge of simultaneously selecting diagnostic questions and evaluating mental health status li2025aligning; sener2018multi. Furthermore, maximizing the diagnostic yield within the constraints of limited clinician-patient interaction time remains a key priority, as excessive probing, especially via LLM-based dialogue agents can be mentally taxing and lead to decision fatigue or disengagement jin2025don, highlighting the need for adaptive systems that select only the most informative and relevant follow-up questions.
@@
To address these challenges, we propose MAQuA: an adaptive, language-based assessment framework that supports multidimensional mental health modeling. This framework infers multiple underlying condition scores while adaptively selecting the most informative follow-up questions, thereby guiding interactions efficiently toward richer diagnostic insight, making it suitable to operate alongside LLM agents. Building on item response theory (IRT)-based adaptive assessments introduced by  varadarajan-etal-2024-alba, our results demonstrate that optimizing information gain across multiple conditions simultaneously can be even more effective than single-score models.
@@
To explore this, we empirically benchmark single-task and multitask models on a multidimensional mental health dataset and assess the effectiveness of adaptive question selection. Our research investigates (1) whether cross-condition information sharing improves per-condition predictive performance; (2) validity of automatically inferred mental health dimensions; and (3) the efficacy of multidimensional IRT in sustaining validity of measures in subsequent adaptive assessment turns. Our main contributions include: (a) a systematic comparison of multi-condition modeling techniques, (b) an adaptive assessment framework for optimizing information gain across outcomes, (c) empirical results demonstrating robust gains in multidimensional assessment, and (d) the release of a novel, questionnaire-driven dataset to support future research.
2 Related Work
@@
While large language models (LLMs) have shown promising results in zero-shot and few-shot prediction tasks (ganesan2024explaining; hur2024language), finetuned or instruction-tuned models remain generally more reliable and better validated across a range of mental health outcomes (xu2024mental). Recent advances demonstrate that open-ended language responses to standardized questions can predict mental health scores with high accuracies: sometimes with correlations exceeding 0.8 with established clinical rating scales (kjell2022natural; varadarajan-etal-2024-alba; sikstrom2023precise). These include pre-trained language models tailored for mental healthcare applications, such as ClinicalBERT and MentalBERT (alsentzer2019publicly; ji2022mentalbert).
@@
Large language models (LLMs) commonly address referential and vague queries by employing targeted, selective prompts, which has been shown to improve answer accuracy and reduce errors zhang2025modeling; kuhn2022clam. However, despite these improvements, LLMs still fall short of human conversational subtlety and adaptiveness when it comes to clarification and follow-up questions. In terms of LLMs for mental health support, rosenman-etal-2024-llm demonstrate that LLMs can effectively transform unstructured psychological interviews into structured questionnaires, enabling automated, multidimensional psychiatric evaluation, though reliability and consistency still require further improvement for clinical deployment. Complementing this, nguyen-etal-2025-large explore LLMs’ ability to engage in mental health counseling, showing that, though LLMs can generate contextually relevant follow-up questions, they often lag behind human clinicians in empathy, specificity, and diagnostic nuance, and in crafting clarifying or probing questions that are crucial for effective counseling. Similarly, yang-etal-2023-towards assess LLM performance across a spectrum of mental health tasks, finding that LLMs frequently overlook emotional cues or oversimplify questions, limiting their utility for nuanced clinical interpretation. This underscores the need for methods like ours that explicitly model multiple mental health factors to guide strategic question selection. Our framework enhances both the efficiency and accuracy of mental health assessments by optimizing inquiry and serving as a comprehensive diagnostic tool.
@@
While adaptive testing has gained significant traction in educational settings, its multidimensional applications remain relatively underexplored, particularly when leveraging open-ended language responses. Most existing approaches in NLP rely on unidimensional item response theory (IRT) focusing on single outcomes (lalor-etal-2016-building; varadarajan-etal-2024-alba). To date, no prior work has effectively bridged this gap by integrating adaptive item selection with multitask learning for language-based, multidimensional mental health evaluation.
@@ 3 Background
Item Response Theory (IRT) is a probabilistic, data-driven measurement framework that models the relationship between an individual’s latent trait score (such as depression severity) and their probability of specific item responses on a questionnaire embretson2000item; hambleton1991fundamentals. In single-dimensional IRT, this relationship is modeled with respect to one latent trait (usually denoted $\theta$), accounting for item-specific parameters such as difficulty and discrimination. IRT enables precise ordering and calibration of items based on their informativeness in measuring the latent trait, supporting adaptive and individualized assessment. Adaptive language-based assessments (varadarajan-etal-2024-alba) were first introduced using single-dimensional IRT, summarized in Algorithm 1.
Exploratory Factor analysis (EFA) is a statistical technique used to uncover the underlying structure of a set of observed variables by identifying clusters of variables that co-vary together, known as factors or latent constructs (cudeck2000exploratory). Although it is very similar to Principal Component Analysis (PCA), it captures the multidimensional, often overlapping nature of latent variables by allowing factors to be correlated, unlike PCA which assumes uncorrelated components. This approach captures meaningful psychological constructs like depression and anxiety, supporting the development of sensitive, multidimensional assessments for accurate mental health assessments. EFA allows us to distill large and complex data such as responses to multiple questionnaires assessing various psychological conditions. Each factor represents a distinct underlying construct that accounts for shared variance among the observed variables, providing insight into how different mental health symptoms or traits may be related at a deeper level.
@@
4 Dataset
To explore these questions, we selected a unique set of material including language questions, diagnoses, and validated clinical scales. The data was collected in two phases: first, we pre-screened the participants to establish a diverse sampling pool including participants who stated to were diagnosed by a mental health professional. We focused on several common mental disorders, such as mood disorders (i.e., major depression, generalized anxiety disorder, bipolar disorder), autism spectrum disorder, attention-deficit/hyperactivity disorder, eating disorders, obsessive-compulsive disorder, post-traumatic stress disorder, and substance use disorders (i.e., alcohol and/or drug abuse). To enhance ecological validity, we intentionally included participants with comorbidities or co-occurring mental health illnesses, as they are the rule rather than the exception in clinical practice. In a second phase, we collected mental health data from these screened individuals using standardized rating scales, free-text narratives, and open-response questions targeting nine common mental disorders.
We recruited a small set of 515 participants (297 female, 182 male, 35 non-binary, 1 who preferred not to categorize their gender identity) who were diagnosed with any of nine mental health conditions (Anxiety Disorders (AD), Bipolar Disorder (BD), Depression (D), Attention Deficit Hyperactivity Disorder (ADHD), Post-Traumatic Stress Disorder (PTSD), Obsessive-Compulsive Disorder (OCD), Eating Disorders (ED), Addiction and Substance Abuse (A), Autism (AU)). The age of the sample ranged from 18 to 78 (mean = 38.9, SD = 12.3) We screened participants for having one (or more) ongoing mental disorders (e.g., Np = 515, $\sim$50 of each disorder). The distribution of all the diagnoses across the participants is shown in Figure A.1. The participants first took a screening questionnaire for diagnoses and treatment for the ten mental health conditions to qualify for eligibility to participate in the survey. Here, 186 individuals reported receiving no treatment at all, while 329 reported receiving at least one form of treatment (i.e., medication and/ or psychotherapy).
@@ 4 Dataset
Ten validated clinical scales were also administered: the PHQ-9 ( kroenke2001phq; 9 items, 4-point Likert) for depression, GAD-7 (spitzer2006brief; 7 items, 4-point Likert) for anxiety, MDQ (miller2004sensitivity; 14 binary items plus one 4-point Likert item) for bipolar disorder, RAADS-14 (eriksson2013raads; 14 items, 4-point Likert) for autism, ASRS Part A (adler2006validity; 6 items, 5-point Likert) for ADHD, NSESSS-PTSD (lebeau2014dimensional; 9 items, 5-point Likert plus an open-text trauma description) for PTSD, BOCS (goodman1989yale; 15 items, 3-point Likert plus an open-response categorization) for obsessive-compulsive symptoms, EDE-QS (fairburn2008eating; 12 items, 4-point Likert) for eating disorders, and two substance use instruments: the AUDIT (allen1997review; 8 items, 5-point Likert plus 2 items, 3-point Likert) for alcohol misuse and the DUDIT (berman2007dudit; 9 items, 5-point Likert plus 2 items, 3-point Likert) for drug abuse. Post-screening, we invited the selected participants to participate in the main study, with recruitment conducted via Prolific. The language questions, some eliciting descriptive words and some open-ended essay-like responses, are listed in Table A.2. They were asked in random order to eliminate any systematic priming effects.
5 Methods
@@
5.1 Multi-outcome Modeling
Given the multitude of psychological dimensions, associated language data, and the occurrence of comorbid diagnoses, linguistic expressions intended to capture one dimension may also provide valuable information about others. To investigate this, we frame the language-based modeling of psychological scores in two configurations: single-task and multi-task. Using a stratified sampling approach based on depression outcomes (PHQ scores) over the users, we generate 9 folds for evaluation, 5 as train set, 3 as development set and 1 test set. Each fold contains the responses of around 58 users each. Language representations are derived from all text responses utilizing an encoder model. We experimented with popular models as shown in Table 4, and for the rest of the experiments, we used the nomic-embed-text-v1.5 model nussbaum2024nomic, which are subsequently reduced to 16 dimensions through Matryoshka embeddings kusupati2022matryoshka. We then train linear regression models to predict standard mental health questionnaire scores across ten dimensions (including PHQ for depression, GAD for anxiety etc. as explained in §4). The single-task models predict each mental health dimension independently, whereas the multi-task models simultaneously predict all ten scores (see Figure 1). For model optimization, we performed hyperparameter tuning over ranges of learning rates, weight decay values, and used the Adam optimizer. We also compared output scaling methods for the regression, finding that min-max scaling consistently outperformed z-score normalization across all settings. Experiments were conducted using a single NVIDIA A6000 GPU. Further, we also prompt llama3.2-1B model grattafiori2024llama to compare the capacilities of zero-shot models in estimating the mental health condition scores (A.4). We also explore several model variants to analyze different aspects of the language data.
Comparison of Embedding models
We compare various encoder models for multi-outcome modeling, the results are shown in Table 4. We find that Matryoshka reduction on the number of dimensions helps with modeling the less represented conditions (across all the conditions other than Depression and Anxiety).
@@ 5.1 Multi-outcome Modeling
7:  Set initial latent trait estimates $\bm{\theta}_{i}^{(0)}$ for each individual $i$
@@
9:while stopping criteria not met on $\bm{\theta}_{i}^{(t)}$ and pool not empty do
10:  For each candidate item $p$, compute Fisher information matrix $\mathcal{I}_{p}(\bm{\theta}_{i}^{(t)})$
11:  Select next item $p^{*}$ maximizing $\det(\mathcal{I}_{p^{*}}(\bm{\theta}_{i}^{(t)}))$ over all candidates (D-optimality)
@@
15:  Update $\bm{\theta}_{i}^{(t+1)}$ using maximum likelihood on responses
@@
18:Output: Final multidimensional trait estimates $\bm{\theta}_{i}^{(t)}$ for individual $i$
@@
Clinical Scales (upper) .404 .423 .440 .454 .097 .133 .073 .172 .080 - .253
LLaMA-3.2 (zero-shot) .032 .104 .036 034 .065 -.005 .029 .071 .069 - .048
@@
1. Aggregation type
2. Question Information
Model
Depre-
ssion
Anxiety Bipolar Autism
Substance
use
OCD ADHD PTSD Eating
Alcohol
use
Avg. Number of dim
nomic-embed-text-v1 .784 .722 .446 .449 .419 .570 .560 .532 .468 .478 .543 16
nomic-embed-text-v1 .763 .688 .442 .487 .239 .426 .403 .265 .270 .232 .421 768
mxbai-embed-large-v1 .758 .668 .480 .484 .246 .434 .406 .240 .298 .150 .417 1024
roberta-base .733 .658 .394 .428 .148 .441 .298 .234 .252 .231 .382 768
roberta-large .780 .675 .385 .449 .210 .429 .418 .269 .274 .225 .411 1024
Table 4: Comparison of popular embeddings nussbaum2024nomic; emb2024mxbai; liu2019roberta derived from encoder models, for multitask, input-aggregated setting. While any of these models could be used with our framework, we find Matryoshka reduction to be particularly helpful in modeling in low-resource settings.
5.2 Factor Analysis
Multidimensional adaptive question-asking algorithm requires setting up a factor model, which is defined through factor analysis to determine the factor structure, that takes in all the questions. We run exploratory factor analysis to determine the optimal factor structure for this dataset for the predicted user-level scores. The factor loadings are reported in Table 3 and Figure A.2. Then we run aggregate the multi-outcome model predictions at a question level, across all users, to analyze how each question loads on to the factors by by applying the factor model at the question-level. This yields a question-level loading which is used to define the MIRT model (Table 5).
Factor Question IDs Condition/Symptom
Table 5: Question-level mapping to the two significant factors that was further used to specify the MIRT model (see Table A.1 for legend).